With the aid of a string, a gyroscope is accelerated from rest to 40 rad/s in 0.40 s under a constant angular acceleration. (a) What is its angular acceleration in rad/s²? (b) How many revolutions does it go through in the process?

a. 16 ra d.ls2 ana 0.2 rev

b. 16 radls2 and 0.4 rev

c. 100 rad/s2 and 0.7 rev

,d. 100 radls2 and 1.3 rev
Answer #1

Apply rotational kinematic equation

wf- wi = alpha t

alpha = 40-0/0.40

=100 rad/s^2

theta = 1/2 * alpha t^2

= 1/2 * 100 (0.4)^2

= 8 rad ( 1 rev/ 2 pi rad) = 1.3 rev

option (d) is correct answer